# Distributed Multiple Fault Detection and Estimation in DC Microgrids with Unknown Power Loads

### v1

@@ 4.4.2. Estimation of aggregated faulty line current
Upon successful detection of any faulty line current $f_{I,i}$ in DG unit $i$, an estimator is activated. The design method of the estimator is provided in this part. We begin by reformulating the parity-space equation (22) based on (20) to account for power line faults and obtain:
@@
Theorem 4.3 (Performance bound of the regularized estimator).
where $\bar{\lambda}_{\mathcal{M}_{i,T}}=\Psi^{\top}_{i,T}\Psi_{i,T}+\bar{\mathcal{Z}}^{\top}_{i,T}\bar{\mathcal{Z}}_{i,T}$ and $\underline{\lambda}_{\mathcal{M}_{i,T}}=\frac{\Psi^{\top}_{i,T}\Psi_{i,T}\bar{\mathcal{Z}}^{\top}_{i,T}\bar{\mathcal{Z}}_{i,T}-\Psi^{\top}_{i,T}\bar{\mathcal{Z}}_{i,T}\bar{\mathcal{Z}}^{\top}_{i,T}\Psi_{i,T}}{\Psi^{\top}_{i,T}\Psi_{i,T}+\bar{\mathcal{Z}}^{\top}_{i,T}\bar{\mathcal{Z}}_{i,T}}$.
Proof.
The proof is relegated to Appendix B. ∎
  1. (1)
    Ill-conditioning issue. When $V_{i}(k)$ approaches a constant value, $\underline{\lambda}_{\mathcal{M}_{i,T}}$ in the denominator of (4.3) tends to zero. This makes the estimation problem ill-conditioned without the regularized term in (32), i.e., $\eta=0$.
  2. (2)
    Variation terms $\Delta\bm{f}_{I_{i},T-1}$ and $\Delta\bm{P}_{i}$. The bound is affected by the fault deviation $\Delta\bm{f}_{I_{i},T-1}$ and the load change $\Delta\bm{P}_{i}$, whose effect can be mitigated effectively by choosing a large enough $\eta$.
  3. (3)
Remark 4.4 (Selection of $\eta$).
The right-hand side of inequalities (4.3) indicates that increasing $\eta$ in (32) can reduce the expectation of the estimation error. Therefore, a large enough $\eta$ needs to be chosen to ensure estimation performance, while being numerically bounded for practical considerations.
Remark 4.5 (Location of the faulty line).
Faults on a power line can be detected and estimated by the diagnosis components of the DG units located at both ends of the line. By analyzing the estimated faulty line currents, the faulty line can be identified in some scenarios. For instance, in the simplest case where only one line in the DC microgrid is faulty, the fault location can be precisely determined.
Remark 4.6 (Conservativeness analysis).
While the regularization term introduced in (32) resolves the issue of the matrix $\Gamma_{i,T}$ being close to singularity, it restricts the variation of the estimate of $P_{i}$ relative to its previous value. Consequently, the proposed estimation approach exhibits limited accuracy in estimating step load changes during the estimation phase. Nevertheless, the estimation performance for faulty line current $f_{I,i}$ improves by increasing the value of $\eta$ as discussed above. Moreover, we observed in the simulation results that the impact of load changes on the estimation of $f_{I,i}$ is limited.
@@ II. Estimation of the faulty line current $f_{I,i}$
    Generate $\tilde{r}_{i}$ using pre-filter 1 in (12), line current estimators in (14), pre-filter 2 in (16), and (20)
@@ 5. Simulation Results
In this section, the effectiveness of the proposed diagnosis scheme is validated on the DC microgrid depicted in Fig. 1, a simplified structure derived from [35]. The parameters of the microgrid are provided in Table 1, with the reference voltages set as $^{*}_{1}=48 V, V$$^{*}_{2}=48.1 V$, and $V^{*}_{3}=47.5$ V. Suppose that the microgrid is initially working under normal conditions. Then, faults might occur when the system is at or close to steady state. The initial conditions are thus set as $x_{1}(0)=[47.80,6.93,11.19]^{\top},x_{2}(0)=[48.10,4.76,15.61]^{\top},\text{and}~x_{3}(0)=[47.50,-4.29,15.02]^{\top}.$ Additionally, since a discretization approach is employed to estimate the faulty current $f_{I,i}$, all simulation results are presented using sampling data to ensure consistency, with a sampling time of $1\times 10^{-5}$ s. Based on the above setup, the diagnosis performance of $\mathcal{D}_{a,i}$ and $\mathcal{D}_{l,i}$ is evaluated separately in the subsequent subsections.
@@ 5.1. Actuator Fault Diagnosis Results
To estimate actuator faults, an estimator $\mathcal{D}_{a,i}$ for each DG unit is designed by utilizing the approach proposed in Section 3. Taking $\mathcal{D}_{a,1}$ as an example, the design process is outlined as follows:
@@
Step 2. Obtain the parameters of the filter’s numerator by solving the optimization problem (10) with the YALMIP toolbox [36].
Step 3. Construct the filter in (6). The estimation of $f_{a,1}$ can be generated by processing the locally available signals $y_{1}$ and $V^{*}_{1}$ through the designed filter.
Suppose an incipient fault $f_{a,1}$ occurs in the voltage controller of DG $1$ and a step fault $f_{a,3}$ occurs in the voltage controller of DG $3$. The incipient fault $f_{a,1}(t)=0$ for $t\leq 100$ ms, and for $t\geq 100$ ms, $f_{a,1}(t)$ evolves as:
@@
where $\beta_{a}=8\times 10^{-9}$ determines the changing rate and stealthiness of $f_{a,1}$ and $\bar{f}_{a,1}=5$ is the final value of $f_{a,1}$. The step fault $f_{a,3}$ remains 0 until $t=70$ ms, after which it becomes $0.5$, i.e., $f_{a,3}(t)=0.5$ for $t>70$ ms.
To demonstrate that the designed estimator can decouple the effects of load changes and power line faults, we consider a step change in $P_{1}$, i.e., $P_{1}(t)=100$ W for $t\leq 40$ ms and $P_{1}(t)=120$ W for $t>40$ ms, as well as the occurrence of a power line fault between DG $1$ and DG $2$, i.e., $f_{L,1}(t)=0$ for $t\leq 80$ ms and $f_{L,1}(t)=0.1$ for $t>80$ ms. The power load $P_{2}$ changes from $125$ W to $110$ W at $120$ ms, and $P_{3}(t)=140$ W is constant. Moreover, noise $\delta_{i}$ and $\zeta_{i}$ for $i=\{1,2,3\}$, $\epsilon_{k}$ for $k=\{1,2\}$ are zero mean white noise with standard deviations $0.01$, $0.001$, and $0.01$, respectively.
The actuator fault estimation results are presented in Fig. 4 and Fig. 5. Specifically, Fig. 4 depicts the voltage and current variations within the microgrid in response to load changes and system faults. As can be seen that the measured signals $V_{i}$ and $I_{t,i}$ do not provide clear indications of fault occurrences. Fig. 5 shows the diagnosis results from each local actuator fault estimator. For brevity, we just analyze the diagnosis results of $\mathcal{D}_{a,1}$ in Fig. 5, as the remaining results follow a similar analysis procedure. The first two subplots of Fig. 5 illustrate the load change in $P_{1}$ and the faulty current $f_{I,1}$, while the third subplot presents the actuator fault $f_{a,1}$ and its estimate $\hat{f}_{a,1}$. It can be observed that the estimated value $\hat{f}_{a,1}$ accurately follows the true fault signal $f_{a,1}$ and remains unaffected by load changes, power line faults, and neighboring DG unit dynamics.
Figure 4. Dynamic response of the DC microgrid when considering $f_{a,1}$, $f_{a,3}$, $f_{L,1}$, and load changes.
Figure 5. Diagnosis of actuator faults: (a) diagnosis results of $\mathcal{D}_{a,1}$, (b) diagnosis results of $\mathcal{D}_{a,2}$, and (c) diagnosis results of $\mathcal{D}_{a,3}$.
@@ 5.2. Power Line Fault Diagnosis Results
For the estimation of faulty line current $f_{I,i}$ induced by power line faults, we introduce the block $\mathcal{D}_{l,i}$ for each DG unit, whose structure is illustrated in Fig. 3. The proposed estimation scheme is compared with the multiple fault estimation method developed in [30] to demonstrate the superiority. Taking $\mathcal{D}_{l,1}$ as an example, its design process is summarized as follows:
Step 1. Design Pre-filter 1 depicted by (12) in Section 4.A. We choose the degree $d_{\mathcal{N}}=2$ and the denominator $a(p)=(0.5+p)(0.1+p)(1+p)$. The coefficients of $\mathcal{N}_{1}(p)$ are determined using the approach outlined in Proposition 3.1. The residual generator (12) takes locally available signals $y_{1}$ and $V^{*}_{1}$ as inputs and produces the output $r_{1}$.
Step 2. Design line current estimators described by (14) in Section 4.B for all power lines connected to DG $1$. These line current estimators use the PCC voltages at the ends of the power lines (i.e., $V_{1},V_{2}$, and $V_{3}$ for DG $1$) as inputs and generate the estimated fault-free component of the line currents $\hat{I}_{1,h}$ and $\hat{I}_{2,h}$.
Step 3. Design Pre-filter $2$ described by (16) in Section 4.C and generate the residual $\tilde{r}_{1}$ used for faulty line current estimation. The input of Pre-filter $2$ is the sum of the estimated healthy line currents $\hat{I}_{1,h}$ and $\hat{I}_{2,h}$. Then, the residual $\tilde{r}_{1}$ can be obtained based on (17).
Step 4. Design the faulty line current estimator for $f_{I,1}$ based on (31)-(33). Discretize the system with a sampling time of $1\times 10^{-5}$ s and choose a sliding window length of $T=20$. The residual $\tilde{r}_{1}$ is first used to distinguish between step load changes and power line faults according to Proposition 4.1. The estimator (33) is activated when a power line fault is detected with the weight $\eta=1\times 10^{6}$, which produces the estimate of $\bar{f}_{I_{1},T-1}$.
@@
Figure 7. Diagnosis of incipient power line faults: (a) diagnosis results of $\mathcal{D}_{l,1}$, (b) diagnosis results of $\mathcal{D}_{l,2}$, and (c) diagnosis results of $\mathcal{D}_{l,3}$.
Figs. 6-8 present the diagnosis results of the incipient power line fault $f_{L,1}$. Particularly, Fig. 6 illustrates the voltage and current variations of each DG unit under load changes and system faults. Due to the stealthiness of the fault, the occurrence of $f_{L,1}$ cannot be discerned directly from these measurements. Fig. 7 depicts the diagnosis result of the faulty line current estimator of each DG unit. For instance, Fig. 7(a) shows the results of $\mathcal{D}_{l,1}$. At $t=40$ ms, the step change in $P_{1}$ is detected according to the diagnosis rules in (29), as the status indicator $\sigma_{1}$ (the third subfigure) briefly switch to $1$ with a duration shorter than the slidng window lenght $T$. At $t=80$ ms, the fault $f_{L,1}$ happens and $\sigma_{1}$ becomes $2$, meaning the successful detection of the power line fault $f_{L,1}$.
The bottom sub-figure of Fig. 7(a) compares the estimated values of the faulty current $f_{I,1}$ induced by $f_{L,1}$ using the proposed approach and that in [30]. As can be seen, the proposed estimation approach tracks the faulty current well, while the method in [30] exhibits significant estimation errors and oscillations due to the near-singularity problem. Note that the actuator fault $f_{a,1}$ and dynamics of other DG units have no effects on the diagnosis results of $\mathcal{D}_{l,1}$. Additionally, the power line fault $f_{L,1}$ is also detected and estimated by $\mathcal{D}_{l,2}$, as shown in Fig. 7(b). Fig. 8 shows the value $|\mathbf{E}[\bar{f}_{I_{1},T-1}-\hat{f}_{I,1}]|$ and its upper bound given in (4.3). As the faulty current gradually converges to a constant, the upper bound of the error also decreases, which aligns well with the theoretical analysis.
Figure 8. Performance bounds derived by Theorem 4.3.
@@
The parameters in the circuit are $R_{k,1}=R_{k,2}=0.025$$\mathrm{\SIUnitSymbolOhm}$\$, $L_{k,1}=L_{k,2}=1~\mu\text{H}$, and $R_{f}=0.01$$\mathrm{\SIUnitSymbolOhm}$\$. The short-circuit fault happens at $t=100$ ms. The dynamics of the incipient fault $f_{L,2}$ follow: $\dot{f}_{L,2}(t)=-\beta_{l}f_{L,2}(t)+\beta_{l}\bar{f}_{L,2}$, with $\beta_{l}=4\times 10^{-9}$, $\bar{f}_{L,2}=1$, and the occurrence time $t=60$ ms. An actuator fault $f_{a,1}$ is considered in DG unit $1$, which is described by a step signal and changes from 0 to $0.6$ at $t=100$ ms. Additionally, to assess the impact of load variations on faulty current estimation, we introduce a step change in $P_{2}$, which decreases from $115$ W to $105$ W at $t=150$ ms.
@@
Figure 10. Diagnosis of short-circuit line faults: (a) diagnosis results of $\mathcal{D}_{l,1}$, (b) diagnosis results of $\mathcal{D}_{l,2}$, and (c) diagnosis results of $\mathcal{D}_{l,3}$.
@@
Figure 12. Residual $\tilde{\Upsilon}_{1,T}$ for isolation of the incipient power line fault $f_{I,1}$.
The simulation results of the second scenario are presented in Figs. 11-13. In particular, Fig. 11 shows the voltage and current variations of the microgrid under load changes and system faults. When the incipient power line fault $f_{L,2}$ happens, its effect on voltages and currents is unobvious from the measurements. In contrast, the short-circuit fault $f_{L,1}$ at $t=100$ ms introduces a low-resistance $R_{f}$, causing sharp voltage and current increases in the microgrid, as shown in Fig. 11. As a result, the short-circuit fault is easily detectable from measurement signals. Nevertheless, the proposed diagnosis approach can provide more detailed information about these faults.
@@
Figure 13. Performance bounds derived by Theorem 4.3.

### v2

@@ 4.4.2. Estimation of aggregated faulty line current
Upon successful detection of any faulty line current $f_{I,i}$ in DG unit $i$, an estimator is activated. The design method of the estimator is provided in this part. We begin by reformulating the parity-space equation (20) based on (4.3) to account for power line faults and obtain:
@@ II. Estimation of the faulty line current $f_{I,i}$
    Generate $\tilde{r}_{i}$ using pre-filter 1 in (12), line current estimators in (14), pre-filter 2 in (16), and (4.3)
@@
Theorem 4.4 (Performance bound of the regularized estimator).
where $\bar{\lambda}_{\mathcal{M}_{i,T}}=\Psi^{\top}_{i,T}\Psi_{i,T}+\bar{\mathcal{Z}}^{\top}_{i,T}\bar{\mathcal{Z}}_{i,T}$ and
$\underline{\lambda}_{\mathcal{M}_{i,T}}=\frac{\Psi^{\top}_{i,T}\Psi_{i,T}\bar{\mathcal{Z}}^{\top}_{i,T}\bar{\mathcal{Z}}_{i,T}-\Psi^{\top}_{i,T}\bar{\mathcal{Z}}_{i,T}\bar{\mathcal{Z}}^{\top}_{i,T}\Psi_{i,T}}{\Psi^{\top}_{i,T}\Psi_{i,T}+\bar{\mathcal{Z}}^{\top}_{i,T}\bar{\mathcal{Z}}_{i,T}}.$
Proof.
The proof is relegated to Appendix B. ∎
  1. (1)
    Ill-conditioning issue. When $V_{i}(k)$ approaches a constant value, $\underline{\lambda}_{\mathcal{M}_{i,T}}$ in the denominator of (4.4) tends to zero, and thus the bound becomes infinite without the regularized term in (30), i.e., $\eta=0$. This makes the estimation problem ill-conditioned.
  2. (2)
    Variation terms $\Delta\bm{f}_{I_{i},T-1}$, $\Delta\bm{P}_{i}$, and the noise level. The tightness of the bound is mainly affected by $\Delta\bm{f}_{I_{i},T-1}$, $\Delta\bm{P}_{i}$, and $\bar{\lambda}_{\Sigma_{\Omega_{i,T}}}$. Particularly, when the faulty current is constant (or converges to a constant) and the load remains unchanged, the bound is tight. In contrast, it becomes conservative when the faulty current continues to evolve or when step load changes occur. The noise level affects the bound through  $\bar{\lambda}_{\Sigma_{\Omega_{i,T}}}$. Therefore, a larger $\eta$ is generally needed to maintain a relatively tight bound when the noise is smaller, and to mitigate the effects of non-zero $\Delta\bm{f}_{I_{i},T-1}$ and $\Delta\bm{P}_{i}$.
  3. (3)
Remark 4.5 (Selection of $\eta$).
The selection of $\eta$ mainly depends on the ill-conditioning of $\Gamma_{i,T}$, the noise level characterized by $\Sigma_{\Omega_{i,T}}$, and the bias–robustness trade-off. In general, a larger $\eta$ is needed when $\Gamma_{i,T}$ is more ill-conditioned or when the data-fitting term is weighted more heavily. Increasing $\eta$ improves robustness to variations in $P_{i}$ and $f_{I_{i}}$, but may introduce bias if the prior load estimate is inaccurate. We show in the simulation that increasing $\eta$ can reduce the estimation error.
Remark 4.6 (Applicability range of faults and possible extension).
Except for the actuator and power line faults, the proposed diagnosis framework also applies to other faults that can be modeled as additive unknown inputs with known structural directions that do not lie in the subspace spanned by the considered disturbances, such as sensor faults and parameter faults. However, the current design does not explicitly isolate these additional fault types. Fault isolation could be achieved by introducing additional diagnosis filters with structured fault sensitivity.
Remark 4.7 (Complexity analysis).
Although the faulty line current estimation algorithm involves multiple computational steps, its complexity is polynomial in the filter order $d_{\mathcal{N}}$, the input dimension $n_{y}$, and the projected residual dimension $n_{\Upsilon}$, and it does not require iterative online optimization. Consequently, the proposed method incurs a relatively low online computational burden. A detailed analysis of the computational complexity is provided in the Appendix. To further demonstrate the real-time feasibility of the proposed algorithm, we also evaluate its average execution time in simulation.
@@ 5. Simulation Results
In this section, the effectiveness of the proposed diagnosis scheme is validated on the DC microgrid depicted in Fig. 1 [24]. All simulations are performed using MATLAB R2022b on a workstation equipped with an Intel i7-13800H processor (2.50 GHz) and 32 GB RAM. The optimization problems involved in the proposed method are solved using the YALMIP toolbox [13]. The parameters of the microgrid are provided in Table 1, with the reference voltages set as $^{*}_{1}=48 V, V$$^{*}_{2}=48.1 V$, and $V^{*}_{3}=47.5$ V. Suppose that the microgrid is initially working under normal conditions. Then, faults might occur when the system is at or close to steady state. The initial conditions are thus set as $x_{1}(0)=[47.80,6.93,11.19]^{\top},x_{2}(0)=[48.10,4.76,15.61]^{\top},\text{and}~x_{3}(0)=[47.50,-4.29,15.02]^{\top}.$ Additionally, since a discretization approach is employed to estimate the faulty current $f_{I,i}$, all simulation results are presented using sampling data to ensure consistency, with a sampling time of $t_{s}=1\times 10^{-5}$ s. Based on the above setup, the diagnosis performance of $\mathcal{D}_{a,i}$ and $\mathcal{D}_{l,i}$ is evaluated separately in the subsequent subsections.
@@ 5.1. Actuator Fault Diagnosis Results
To estimate actuator faults, an estimator $\mathcal{D}_{a,i}$ for each DG unit is designed by utilizing the approach proposed in Section 3. To demonstrate the advantages of the estimation filter constructed within the DAE framework, we compare it with a sliding mode observer (SMO) [8] and an adaptive observer (AO) [34]. Taking $\mathcal{D}_{a,1}$ as an example, the design process is outlined as follows:
@@
Step 2. Obtain the parameters of the filter’s numerator by solving the optimization problem (10) and construct the filter in (6).
Suppose that an incipient fault $f_{a,1}$ occurs in the voltage controller of DG $1$ and a step fault $f_{a,3}$ occurs in the voltage controller of DG $3$. The incipient fault $f_{a,1}(t)=0$ for $t\leq 100$ ms, and for $t\geq 100$ ms, $f_{a,1}(t)$ evolves as:
@@
where $\beta_{a}=8\times 10^{-9}$ determines the changing rate and stealthiness of $f_{a,1}$ and $\bar{f}_{a,1}=5$ is the final value of $f_{a,1}$. The step fault $f_{a,3}$ remains 0 until $t=70$ ms, after which it becomes $2$.
To demonstrate that the designed estimator can decouple the effects of load changes and power line faults, we consider a step change in $P_{1}$, i.e., $P_{1}$ steps from $100$ to $120$ at $40$ ms, as well as the occurrence of a power line fault between DG $1$ and DG $2$, i.e., $f_{L,1}$ steps from 0 to 0.1 at 80 ms. The power load $P_{2}$ changes from $125$ W to $110$ W at $120$ ms, and $P_{3}(t)=130$ W is constant. Moreover, noise $\delta_{i}$ and $\zeta_{i}$ for $i=\{1,2,3\}$, $\epsilon_{k}$ for $k=\{1,2\}$ are zero mean white noise with standard deviations $0.01$, $0.001$, and $0.01$, respectively.
The actuator fault estimation results are presented in Fig. 4 and Fig. 5. Specifically, Fig. 4 depicts the voltage and current in response to load changes and faults. As shown, the measured signals $V_{i}$ and $I_{t,i}$ do not provide clear indications of fault occurrence. Fig. 5 presents the diagnosis results from the local actuator fault estimators. For brevity, we focus on the results of $\mathcal{D}_{a,1}$ depicted in Fig. 5 (a), where the actuator fault $f_{a,1}$ and its estimates generated by (6), AO, and SMO are compared. The results show that the proposed estimate $\hat{f}_{a,1}$ accurately follows the true fault signal $f_{a,1}$ and remains unaffected by load changes, power line faults, and neighboring DG unit dynamics. By contrast, both AO and SMO methods respond more slowly after the fault occurs, while the SMO-based approach also exhibits unavoidable chattering in the estimation results.
Figure 4. Dynamic response of the DC microgrid when considering the actuator fault.
Figure 5. Diagnosis of the actuator fault: (a) results of $\mathcal{D}_{a,1}$, (b) results of $\mathcal{D}_{a,2}$, and (c) results of $\mathcal{D}_{a,3}$.
@@ 5.2. Power Line Fault Diagnosis Results
For the estimation of faulty line current $f_{I,i}$ induced by power line faults, we introduce the block $\mathcal{D}_{l,i}$ for each DG unit, whose structure is illustrated in Fig. 3. The proposed estimation scheme is compared with the multiple fault estimation method developed in [32]. Moreover, to better capture the behavior of practical microgrid systems, we construct a high-fidelity microgrid model in Simulink/Simscape Electrical using electrical and power electronics components with dynamic characteristics. The proposed diagnosis method is then validated on this model to demonstrate its potential practical applicability. Taking $\mathcal{D}_{l,1}$ as an example, its design process is summarized as follows:
Step 1. Design Pre-filter 1 depicted by (12) in Section 4.1. We choose the degree $d_{\mathcal{N}}=2$ and the denominator $a(p)=(0.5+p)(0.1+p)(1+p)$. The coefficients of $\mathcal{N}_{1}(p)$ are determined using the approach outlined in Proposition 3.1.
Step 2. Design line current estimators described by (14) in Section 4.2 for all power lines connected to DG $1$.
Step 3. Design Pre-filter $2$ described by (16) in Section 4.3 and generate the residual $\tilde{r}_{1}$ used for faulty line current estimation.
Step 4. Design the faulty line current estimator for $f_{I,1}$ based on (29)-(31). Discretize the system with a sampling time of $t_{s}=1\times 10^{-5}$ s and choose a sliding window length of $T=20$. The estimator (31) is activated when a power line fault is detected, which produces the estimate of $\bar{f}_{I_{1},T-1}$.
@@
Figure 7. Diagnosis of the incipient power line fault: (a) results of $\mathcal{D}_{l,1}$, (b) results of $\mathcal{D}_{l,2}$, and (c) results of $\mathcal{D}_{l,3}$.
Figs. 6-8 present the diagnosis results of the incipient power line fault $f_{L,1}$. Particularly, Fig. 6 illustrates the voltage and current variations of each DG unit under load changes and faults. Due to the stealthiness of the fault, the occurrence of $f_{L,1}$ cannot be discerned directly from these measurements. Fig. 7 depicts the diagnosis results of the faulty line current estimator of each DG unit. For instance, Fig. 7(a) shows the results of $\mathcal{D}_{l,1}$. At $t=40$ ms, the step change in $P_{1}$ is detected according to the diagnosis rules in (27), as the status indicator $\sigma_{1}$ briefly switches to $1$ with a duration shorter than $T$. At $t=80$ ms, the fault $f_{L,1}$ happens and $\sigma_{1}$ becomes $2$, meaning the successful detection of the power line fault $f_{L,1}$.
The bottom sub-figure of Fig. 7(a) compares the estimates of the faulty current $f_{I,1}$ obtained using the proposed DAE-based approach, the Simulink model, and the estimation approach in [32]. As shown, the proposed estimation approach tracks the faulty current well, whereas the method in [32] exhibits significant estimation errors and oscillations due to near-singularity. The Simulink-based results are consistent with those obtained from the MATLAB implementation, further supporting the practical applicability of the proposed diagnosis framework. It is also worth noting that the actuator fault $f_{a,1}$ and dynamics of other DG units have no effects on the diagnosis results of $\mathcal{D}_{l,1}$. Additionally, the power line fault $f_{L,1}$ is also detected and estimated by $\mathcal{D}_{l,2}$, as shown in Fig. 7 (b). Fig. 8 presents the value $|\mathbf{E}[\bar{f}_{I_{1},T-1}-\hat{f}_{I,1}]|$ and its upper bound given in (4.4). As the faulty current gradually converges to a constant, the upper bound of the error also decreases, which aligns well with the theoretical analysis. Fig. 9 further illustrates the effect of increasing $\eta$ in reducing the estimation error.
Figure 8. Performance bounds derived by Theorem 4.4 under the incipient power line fault.
Figure 9. Expectation of the fault line current estimation error under different $\eta$.
@@
The parameters in the circuit are $R_{k,1}=R_{k,2}=0.025$$\mathrm{\SIUnitSymbolOhm}$\$, $L_{k,1}=L_{k,2}=1~\mu\text{H}$, and $R_{f}=0.01$$\mathrm{\SIUnitSymbolOhm}$\$. The short-circuit fault happens at $t=100$ ms. The dynamics of the incipient fault $f_{L,2}$ follow: $\dot{f}_{L,2}(t)=-\beta_{l}f_{L,2}(t)+\beta_{l}\bar{f}_{L,2}$, with $\beta_{l}=4\times 10^{-9}$, $\bar{f}_{L,2}=1$, and the occurrence time $t=50$ ms. An actuator fault $f_{a,1}$ is considered in DG unit $1$, which is described by a step signal and changes from 0 to $0.6$ at $t=100$ ms. Additionally, to assess the impact of load variations on faulty current estimation, we introduce a step change in $P_{2}$, which decreases from $115$ W to $105$ W at $t=150$ ms.
@@
Figure 12. Diagnosis of the incipient and short-circuit line faults: (a) results of $\mathcal{D}_{l,1}$, (b) results of $\mathcal{D}_{l,2}$, and (c) results of $\mathcal{D}_{l,3}$.
Figure 13. Residual $\tilde{\Upsilon}_{1,T}$ for isolation of the incipient and short-circuit line faults.
The simulation results of the second scenario are presented in Figs. 11-14. In particular, Fig. 11 shows the voltage and current variations of the microgrid under load changes and system faults. When the incipient power line fault $f_{L,2}$ happens, its effect on voltages and currents is unobvious from the measurements. In contrast, the short-circuit fault $f_{L,1}$ at $t=100$ ms introduces a low-resistance $R_{f}$, causing sharp voltage and current increases in the microgrid, as shown in Fig. 11. As a result, the short-circuit fault is easily detectable from measurement signals. Nevertheless, the proposed diagnosis approach can provide more detailed information about these faults.
@@
Figure 14. Performance bounds derived by Theorem 4.4 under the incipient and short-circuit power line faults.
5.3. Real-time Feasibility Analysis
To evaluate the real-time feasibility of the proposed algorithm, we measured the execution time of one sampling step. The algorithm was executed for $2\times 10^{4}$ iterations on the above computational platform. The average execution time per step is $7.8941\times 10^{-6}$ s, which is smaller than the sampling time $t_{s}=1\times 10^{-5}$ s. This result indicates that the proposed algorithm can be executed within one sampling period and is therefore suitable for real-time implementation. It is worth noting that the reported runtime is obtained from a straightforward MATLAB implementation without dedicated code-level optimization. Further reduction in execution time is expected when the algorithm is implemented using optimized or compiled code.